RANKS OF ELLIPTIC CURVES IN FAMILIES OF QUADRATIC TWISTS

KARL RUBIN AND ALICE SILVERBERG

Abstract. We show that the unboundedness of the ranks of the quadratic twists of an elliptic curve is equivalent to the divergence of certain infinite series.